HUBBLE SPACE TELESCOPE DAILY REPORT #924 PERIOD COVERED: 0800 EST Daily Status Report as of 189/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WFC 4382 (HST Imagery and Spectroscopy of the Dusty SMC H II Region N88A: Cycle 3 High) The Wide Field Camera was used to make six images of the dusty H II region in N88A which is located in the Large Magellanic Cloud. This imagery is to prepare for later spectroscopy, which will be defined in a separate proposal. All six observations were received and displayed. They all were reported to be okay. A recentering event occurred during the last observation. 1.2 Completed Three Sets of FOC/96 4804 (UV Imaging of Nearby Galaxies with FOC/96) The Faint Object Camera was used to complete three ultraviolet snapshots of the nearby galaxy ESO-0324-2. The exposures lasted for ten minutes and used filter F220W. All three observations were received and displayed in the OSS. No anomalies were noted, except for a recentering event which occurred on the last observation. 1.3 Completed Two Sets of FOS/BL 4953 (QSO Absorption System Snapshot Survey Part 3 of 3: Cycle 3 High) The Faint Object Spectrograph (Blue Side) was used to make four observations of quasars 2308+0951 and 1543+4855 to look for absorption lines at low redshift, and also emission lines in the QSO's continuum. All four observations were received and displayed. They were reported to look okay, with maximum counts in the first and third groups of the 1X3 peakups. There was a recentering event which occurred during the last observation. 1.4 Completed FOS/BL 4952 (QSO Absorption System Snapshot Survey - Part 2 of 3 : Cycle 3 High) The Faint Object Spectrograph (Blue Side) was used to make two observations of quasar 1123+5929 to look for absorption lines at low redshift, and also emission lines in the QSO's continuum. Both observations were received and displayed. The 1X3 peakup had maximum counts in the second group, and the spectrum from the following observations looked okay. 1.5 Completed FOS/RD 4396 (QSO Absorption System Snapshot Survey - Part 2 of 3 : Cycle 3 High) The Faint Object Spectrograph (Red Side) was used to make one spectrographic observation of quasar 0506-6113 to look for absorption line systems at low red shift, emission lines, and to study the QSO's continuum. The observations was received in real time. The maximum counts for the 1X3 peakup were in the second group. 1.6 Completed GHRS/2 4966 (Spatially Resolved Spectroscopy of Jupiter - Cycle 3 High) The High Resolution Spectrograph was used to make ultraviolet spectral observations of the planet Jupiter at a given latitude. Blind pointing was used. The planetary rotation was not taken into account, therefore, the resulting spectra will be a zonal average. The observations were executed as scheduled, and no problems were reported. 1.7 Completed WFC 5016 (The Evolution of the UV Spectra in Early Type Galaxies Out to Z = 0.7) The Wide Field Camera was used to make images of the weak radio galaxy 53W039 to study its current morphology. This study will be later followed by proposal 3545 to do a complete spectrographic study of the galaxy. The OSS monitored the observations and assessed the data quality. No problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 11 Successful : 11 Scheduled Reacquisitions : 3 Successful : 3 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 22 Successful : 22 2.3 Significant HSTARs: None 2.4 Operations Note: The Rendezvous and Grapple Simulation is in process. This is the second day of a two day simulation. It is scheduled to end this afternoon. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete WFC 5016 (The Evolution of the UV Spectra in Early Type Galaxies Out to Z = 0.7) 3.2 Complete FOS/RD 4952 (QSO Absorption System Snapshot Survey - Part 2 of 3: Cycle 3 High) 3.3 Complete PC 3603 (Snapshot of Protoplanetary Nebulae) 3.4 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.5 Complete FOS/RD 4953 (QSO Absorption System Snapshot Survey - Part 3 of 3: Cycle 3 High) 3.6 Complete PC 5014 (WF/PC Cycle 3 Calibration: PC Flats, Narrow 2) 3.7 Complete FOC/96 4113 (Auroral Imaging of Jupiter With the FOC Cycle 2, February- March 1993) 3.8 Complete FOS/BL 4396 (QSO Absorption System Snapshot Survey - Part 1 of 3: Cycle 3 High) 3.9 Complete FOS/BL 3618 (Excitation Processes for the Outer Planet UV Emissions: Future - Cycle C) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time.